OpenStudy (anonymous):

find the average values of function over the corresponding intervals 1/x on [1,4]

OpenStudy (anonymous):

please ????

OpenStudy (tkhunny):

\(\dfrac{\int\limits_{1}^{4}\dfrac{1}{x}\; dx}{4-1}\) Are we ringing bells? Think about a rectangle, but shred the top a little.

OpenStudy (anonymous):

i don't know please tell me

OpenStudy (anonymous):

average value is integral divided by the length of the path what @tkhunny said \[\frac{1}{3}\int_1^4\frac{1}{x}dx\]

OpenStudy (anonymous):

anti derivative of \(\frac{1}{x}\) is \(\ln(x)\) so you get \[\frac{1}{3}\left(\ln(4)-\ln(1)\right)\]

OpenStudy (anonymous):

i know the formula for the average but i don't know who i use it in ihis equation

OpenStudy (anonymous):

since \(\ln(1)=0\) this is \(\frac{1}{3}\ln(4)\)

OpenStudy (anonymous):

1/3 write

OpenStudy (anonymous):

please must to solve it by the law (1/b-a)integration fx dx

OpenStudy (anonymous):

so how i start to solve

OpenStudy (anonymous):

i wrote the complete solution above

OpenStudy (anonymous):

thank you